OpenStudy (anonymous):

15x2y3 and 18xy2 A. 3xy B. 90x2y3 C. 30xy2 i feel like its B but the variables dont match up or do they?

OpenStudy (anonymous):

The question does not really make sense!

OpenStudy (anonymous):

What is the least common multiple, LCM, of the two monomials? 15x2y3 and 18xy2 A. 3xy B. 90x2y3 C. 30xy2

OpenStudy (anonymous):

there thats the exact question ^^

OpenStudy (anonymous):

Haha.. Now, that makes sense :)

OpenStudy (anonymous):

^_^

OpenStudy (anonymous):

15x^2y^3=3*5*x^2*y^3 18xy^2=2*3^2*x*y^2

OpenStudy (anonymous):

The LCM=2*3^2*5*x^2*y^3=90x^2y^3

OpenStudy (anonymous):

Hmmm haha i was right then!! ^^ i just didnt do it the same way ahaha thanks anwar! (:

OpenStudy (anonymous):

You're welcome :)